The sound level measured in a room by a person watching a movie on a home theater system varies from
60 dB
during a quiet part to
90 dB
during a loud part. Approximately how many times louder is the latter sound?
1
Expert's answer
2020-12-29T15:29:02-0500

By definition, dB are:



"60dB = 10\lg I_1\\
90dB = 10\lg I_2"

where "I_1, I_2" are intensities of the sounds. From this obtain:


"I_1 = 10^{6}\space (relative\space units)\\
I_2 = 10^{9}\space (relativ\space units)"

Exact units don't matter since we are only interested in ratio "I_2\/I_1". Thus, the latter sound is


"\dfrac{I_2}{I_1} = \dfrac{10^9}{10^6} = 1000"

times louder.


Answer. 1000 times.